A sample of ozone gas contains 7.25 * 10^24 atoms od oxygen-how many molecules of ozone gas are there


1
Expert's answer
2021-04-22T03:04:35-0400

22.04.21

You

09:31

Solution:sample of ozone gas contains 
7.25 * 10^24
 atoms od oxygen-how many molecules 
of ozone gas are there
Answer: N(O)- number of atoms ozone;
n(O)-mole of ozone
N(O)=n(O)*Na that Na is number 
Avogadro ;
n(O)=N(O)/Na=7.25*10^24/6.02*10^23=
=1.2*10=12 mole

 A molecule made up by 3 oxygen atoms
n(O3)=12 mole/3=4 mole 
N(number of molecules)=n*N(A)
N= 4 моль*6,02*10²³ molecules/mole = 
2,408*10²⁴ molecules
Answer: 2,408×10²⁴ molecules